OpenStudy (anonymous):

A + 8 = -32 it says to find the variable?

OpenStudy (ineedhelpnow):

solve for A. subtract 8 from both sides.

OpenStudy (ineedhelpnow):

A+8 -8=-32-8 A=-40

OpenStudy (anonymous):

-40